Find the H.M. of 1/2,2/3,3/4,.......n/(n+1) occurring with frequencies 1,2,3,.....n
1
Expert's answer
2019-10-21T10:42:05-0400

The harmonic mean can be found as


"HM = {{\sum\limits_{k = 1}^n {{f_k}} } \over {\sum\limits_{k = 1}^n {{{{f_k}} \over {{x_k}}}} }}"

where "{{x_k}}" are dara values and "{{f_k}}" are corresponding frequencies. We've got "{f_k} = k" and "{x_k} = {k \over {k + 1}}" . Thus we have to find two sums. The first is just the arithmetic progression


"\sum\limits_{k = 1}^n {{f_k}} = \sum\limits_{k = 1}^n k = {{n(n + 1)} \over 2}"

The second can be simplifid to the arithmetic progression and the sum of units


"\sum\limits_{k = 1}^n {{{{f_k}} \over {{x_k}}}} = \sum\limits_{k = 1}^n {{k \over {{k \over {k + 1}}}}} = \sum\limits_{k = 1}^n {(k + 1)} = \sum\limits_{k = 1}^n k + \sum\limits_{k = 1}^n 1 = {{n(n + 1)} \over 2} + n"

Then


"HM = {{{{n(n + 1)} \over 2}} \over {{{n(n + 1)} \over 2} + n}} = {2 \over {n(n + 1) + 2n}} \cdot {{n(n + 1)} \over 2} = {{n + 1} \over {n + 3}}"